# Theory of Mixture-of-Experts for Mobile Edge Computing

### v1

@@ IV Theoretical Guarantee for Learning Convergence of our MoE Design
$M_{th}=\min\Big{\{}d_{u},\log_{N}\Big{(}\frac{1}{\delta}\Big{)}\Big{\}}$ (16)
with $\delta=o(1)$, for any task $t$ with $\bm{w}_{t}\in\mathcal{W}_{n}$ arrives after the system convergence, with probability at least $1-o(1)$, there always exists an idle expert $m\in\mathcal{M}_{n}$ with $\gamma_{t}^{(m)}=1$ in (2) to provide the correct type of expertise for that task.
@@
Proposition 2 (Routers’ Convergence).
@@ VII Conclusions
We first prove that $|\mathcal{M}_{n}|\geq M_{th}$ guarantees an idle expert for any task arrival with $\bm{w}_{t}\in\mathcal{W}_{n}$ by considering the worst case that tasks within the same cluster consequently arrive for $M_{th}$ time slots.
If $M_{th}=d_{u}$ given the maximum time delay for an MEC node satisfying $d_{u}\leq\log_{N}(\frac{1}{\delta})$, then even if $M_{th}$ task arrivals are in the same cluster, there will be at least one expert that has completed its previous task for the $M_{th}+1$-th task arrival. Therefore, $M_{th}=d_{u}$ guarantees at least an idle expert for any task arrival with $\bm{w}_{t}\in\mathcal{W}_{n}$.
If $M_{th}=\log_{N}(\frac{1}{\delta})$ given $d_{u}>\log_{N}(\frac{1}{\delta})$, the probability that $M_{th}$ consequent task arrivals are within the same cluster is:
$\mathbb{P}(\bm{w}_{d}\in\mathcal{W}_{n},\forall d\in\{t-M_{th},\cdots,t\})=(\frac{1}{N})^{M_{th}}=\delta,$
where the first equality is because that each task arrival has an identical probability of belonging to each cluster, and the second equality is because of $M_{th}=\log_{N}(\frac{1}{\delta})$. Given $\delta=o(1)$, the probability that $M_{th}$ consequent task arrivals are within the same cluster approaches zero. Consequently, $M_{th}=\log_{N}(\frac{1}{\delta})$ guarantees an idle expert for any task arrival with $\bm{w}_{t}\in\mathcal{W}_{n}$ with probability $1-o(1)$.
@@
solving which we obtain $t\geq\lceil M\ln(\frac{M}{\delta})\rceil$. Therefore, given $T_{1}=d_{u}+\lceil\eta^{-1}\sigma_{0}^{-0.5}M\ln(\frac{M}{\delta})\rceil$, each expert has been explored at least once during the expert-exploration stage (i.e., $t<T_{1}$). Here the first inequality is because the probability of $Y_{t}^{(m)}\geq 1$ under routing strategy (4) is greater than that under randomly routing strategy with identical probability $\frac{1}{M}$. The second inequality is because $(1-\frac{1}{M})^{t}\rightarrow 0$ and the third inequality is derived by Binomial theorem.

### v2

@@ IV Theoretical Guarantee for Learning Convergence of our MoE Design
$M_{th}=\frac{d_{u}}{N}+\Phi^{-1}(1-\delta)\sqrt{\frac{d_{u}}{N}(1-\frac{1}{N})}$ (16)
with the standard normal quantile $\Phi^{-1}(\cdot)$ and $\delta=o(1)$, for any task $t$ with $\bm{w}_{t}\in\mathcal{W}_{n}$ arrives after the system convergence, with probability at least $1-o(1)$, there always exists an idle expert $m\in\mathcal{M}_{n}$ with $\gamma_{t}^{(m)}=1$ in (2) to provide the correct type of expertise for that task.
@@
Proposition 2 (Router’s Convergence).
@@ VII Conclusions
We first prove that $|\mathcal{M}_{n}|\geq M_{th}$ guarantees at least an idle expert for any task arrival with $\bm{w}_{t}\in\mathcal{W}_{n}$. For ease of exposition, we named tasks as task $n$ for any task with $\bm{w}_{t}\in\mathcal{W}_{n}$ in the following. In this case, we need to guarantee that the probability, of which task $n$ arrives more than $M_{th}$ times within $d_{u}$ time slots, is smaller than infinitesimal $\delta$. In other words, $M_{th}$ should ensure
[…]
Let $A_{n}$ denote the number of times task $n$ arrives in $d_{u}$ time slots. Since each task is selected independently with probability $\frac{1}{B}$, we have $A_{n}\sim\text{Bionomial}(d,1/N)$, and we seek $\text{Pr}(A_{n}>M_{th})<\delta$.
Since $d_{u}$ is large, we can approximate $A_{n}$ by a normal distribution:
$A_{n}\approx\mathcal{N}(\frac{1}{N},\frac{1}{N}(1-\frac{1}{N})).$
Using the normal tail bound, we have
$\text{Pr}(A_{n}>M_{th})\approx\text{Pr}\left(\frac{A_{n}-d_{u}/N}{\sqrt{d/N(1-1/N)}}>\frac{A_{n}-M_{th}/N}{\sqrt{d/N(1-1/N)}}\right).$
Using the standard normal quantile $\Phi^{-1}(1-\delta)$, we obtain
$M_{th}\geq\frac{d_{u}}{N}+\Phi^{-1}(1-\delta)\sqrt{\frac{d}{N}(1-\frac{1}{N})},$
which guarantees at least an idle expert for task $n$ arrival during any $d_{u}$ time slots with probability $1-\delta$.
@@
solving which we obtain $t\geq\lceil M\ln(\frac{M}{\delta})\rceil$. Therefore, given $T_{1}=d_{u}+\lceil\eta^{-1}\sigma_{0}^{-0.5}M\ln(\frac{M}{\delta})\rceil$, each expert has been explored at least $\eta^{-1}\sigma_{0}^{-0.5}$ times during the expert-exploration stage (i.e., $t<T_{1}$). Here the first inequality is because the probability of $Y_{t}^{(m)}\geq 1$ under routing strategy (4) is greater than that under randomly routing strategy with identical probability $\frac{1}{M}$. The second inequality is because $(1-\frac{1}{M})^{t}\rightarrow 0$ and the third inequality is derived by Binomial theorem.